How do you solve:
-44 + (-3/8) + 95 + (-5/8)

Answer by wuwei96815(245)

-44+(-3/8)+95+(-5/8)
First, remove the parentheses. You get:
-44-3/8+95-5/8
Next, combine like terms. You get:
51-8/8 = 51-1 = 50